• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Yi Qiang

    Log entry time 21:06:03 on February 10, 2009

    Entry number 260371

    keyword=Summary of target optics/cell work since d2n

    Feb 5 swing - Feb 6 owl: Installed two more Q-wave plates for
    longitudinal pumping, EPR fiber change and pickup coils adjustment. Then
    started longitudinal pumping to high energy state.
    Feb 6 day: EPR calibration of longitudinal pumping revealed very low
    polarization ~ 30% after 12 hours pumping
    Feb 6 swing: Re-adjusted the Q-wave plates for longitudinal pumping
    Feb 7 day: Another EPR calibration of longitudinal pumping
    Feb 7 swing - Feb 8 day: 3He in beam with longidtudinal pumping: Maximum
    polarization ~ 47% ( probably 50% w/o beam )
    Feb 8 swing: Rotated the field to transverse, started pumping to high
    energy state
    Feb 9 swing: EPR calibration of transverse pumping, maximum polarization
    ~ 54%, suspect oven temperature then increased the oven temperature from
    230C to 240C. After about two hours, polarization is still about 55%.
    Feb 9 swing - Feb 10 owl: Switch back to longitudinal pumping, average
    polarization ~ 49%.
    Feb 10 day: Change to transverse pumping to low energy state, trying to
    see whether the low polarization was due to masing effect. Until 5:00pm,
    the polarization reached 55% and the projection of maximum shows ~ 60%
    Feb 10 swing: Installed rotatable Q-wave plates to longitudinal line.

    A copy of this log entry has been emailed to: jpchen, moffit, brads